snowflake0531:

If sin Θ = 2 over 3 and tan Θ < 0, what is the value of cos Θ?

snowflake0531:

AZ:

Hey! Do you know in which quadrant \(\sin\theta\) will be positive and \(\tan\theta\) will be negative?

snowflake0531:

yes, sin theta is positive when the y coordinate is positive, and tan is negative when either sin or cos is negative. i'm just thinking either the second or the last one is right

AZ:

That's good! And yes! We're able to eliminate those two answers because we know that cos has to be negative.

1 attachment
snowflake0531:

um so is it the second or the fourth... both has values of a negative number

AZ:

Have you learned about the trigonometric identity \(\sin ^2\theta + \cos^2 \theta = 1\)

snowflake0531:

no? what's that

AZ:

It's called a trigonometric identity but basically it's like equations that involve trig functions that are always true

AZ:

And so we can use that to solve for \(\cos \theta\)

snowflake0531:

uh, how

AZ:

we know that \(\sin\theta = \dfrac{2}{3}\) \(\sin^2\theta = \left(\dfrac{2}{3} \right)^2\) What is \(\left(\dfrac{2}{3} \right)^2 = ?\)

snowflake0531:

4/9

AZ:

\(\sin ^2\theta + \cos^2 \theta = 1\) So now we get \(\dfrac{4}{9} + \cos^2 \theta = 1\) What is 1 - 4/9? And let's try to keep it as a fraction

snowflake0531:

5/9

AZ:

Good! So now we get \(\cos^2\theta = \dfrac{5}{9}\) and that's \(\cos^2 \theta\) and we want \(\cos\theta\) so we take the square root on both sides! Does that make sense? So \(\cos\theta = \sqrt{\dfrac{5}{9}}\) Can you simplify that? Remember \(\sqrt{\dfrac{a}{b}} = \dfrac{\sqrt{a}}{\sqrt{b}}\)

snowflake0531:

it would be sqrt5/sqrt9, simplify- sqrt5/3

snowflake0531:

oh

snowflake0531:

ohhhhhhhh, i see, thxxxxxxxxxxxx

AZ:

Good! So we know it's sqrt(5)/3 But remember the first thing we mentioned? \(\cos\theta\) has to be negative so it'll be -sqrt(5)/3

snowflake0531:

oh, right lol, well, thxx

AZ:

Of course!!! It was my pleasure :)

snowflake0531:

lol

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Good! So we know it's sqrt(5)/3 But remember the first thing we mentioned? \(\cos\theta\) has to be negative so it'll be -sqrt(5)/3 \(\color{#0cbb34}{\text{End of Quote}}\) the - got stuck in the previous line but \( - \dfrac{\sqrt{5}}{3}\)

snowflake0531:

thx

AZ:

You're welcome!